OpenStudy (anonymous):

Find a solution of the differential equation dy/dx=3y+1 that satisfies the initial condition y(0)=2 I got the answer : y=(7e^(-x)-1)/(3) . Am I right?

OpenStudy (turingtest):

yes

OpenStudy (turingtest):

oh no, you forgot the 3 in the exponent

OpenStudy (anonymous):

this is what I get\[y = \frac{7e^{3x} - 1}{3}\]

OpenStudy (turingtest):

right :)

OpenStudy (anonymous):

ops.. can you show me how does it go?

OpenStudy (turingtest):

what did you use as an integrating factor?

OpenStudy (anonymous):

separate the variables\[\frac{dy}{3y + 1} = dx\] integrate both sides\[\frac{1}{3}\ln (3y + 1) = x + C\]solve for y\[y = \frac{e^{3x + 3C} - 1}{3}\]rewrite the constant\[y = \frac{De^{3x}-1}{3}\]apply the initial condition and solve for D

OpenStudy (turingtest):

oh I didn't even notice it was separable I'm tired, g'night :)

OpenStudy (anonymous):

good night and thank both of you guys~ Im finding my mistake now 0.0